I need help solving this system of equations please.
2x+y-3z=7
x-y-2z=4
3x+3y+2z=4

Answer by lynnlo(4176)

(1)x= -y/2+3z/2+7/2
(2)x=y+2z+4
(3)x=-y+4/3-2z/3